graphing by equations linears
x+y=4
x-y=2

Answer by mananth(16946)

x+y=4
x=0 y=4 (0,4)
y=0 x=4 (4,0)
x-y= 2
x=0,y=-2 (0,-2)
y=0x=2 (2,0)

The solution is (3,1)